OpenStudy (anonymous):

Hanna is going on a 8-day hike of 82.5 miles . she plans to hike 8.25 hours everyday for each of those 8 days . how many miles will she cover each hour of her trip ?

OpenStudy (anonymous):

@ccswims can you help me out again please ?

OpenStudy (anonymous):

First, find how many hours she's going to be spending. \[\large \sf Hours~per~day \times Number~of~days~=~Total~number~of~hours\] And imagine there are spaces between the words because latex is acting up /:

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

In what way?

OpenStudy (anonymous):

You said to find out how many hours she'll be spending how do I figure that out ?

OpenStudy (anonymous):

\[\large \sf Hours~per~day \times Number~of~days~=~Total~number~of~hours\]

OpenStudy (anonymous):

How many days are there? How many hours is she hiking per day?

OpenStudy (anonymous):

Its going to be 8 days & she will be hiking 8.25 hours during those 8 days . so your saying I should multiply 8.25× 8 ?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Okay I'll do that

OpenStudy (anonymous):

Okay I multiplied it and got 66.00

OpenStudy (anonymous):

Ok, good. Now all that's left is to find the miles per hour (what the question is looking for). We can find this with the formula \[\bf Number~of~hours \times Miles~per~hour~=~Number~of~miles\]

OpenStudy (anonymous):

Wait I have a question what does the 66 that I multiplied & got represents ??

OpenStudy (anonymous):

The total number of hours they are hiking

OpenStudy (anonymous):

So it should be 66 times 82.5 then ????

OpenStudy (anonymous):

We're looking for the Miles per hour. We already know what the number of miles is.

OpenStudy (anonymous):

Yeah it's 66 so what should I multiply it by ?

OpenStudy (anonymous):

Or should it be 8.25 times 66 ?

OpenStudy (anonymous):

Are u there ?

OpenStudy (anonymous):

I was making dinner sorry. I am back now @Halia . \[\bf Number~of~hours \times Miles~per~hour~=~Number~of~miles\] can be converted to \[\bf Miles~per~hour~=~\frac{Number~of~miles}{Number~of~hours}\]

OpenStudy (anonymous):

Can we finish this off tomorrow ? I'm so exhausted and tired of doing math @legendarySadist